If Regi took out $25129 student loan at a 8% APR., how much interest will he be charged after a year? Answer: Check
Answer #1

Loan Amount = P = $25129

APR = r = 8%

Tenure of Loan = t = 1 year

Amount to be repaid after 1 year = P(1+r)t = 25129(1+0.08)1 = 27139.32

Hence, Interest Paid = Amount Repaid - Loan Amount = 27139.32 - 25129 = $2010.32